calculate de osmotic peessure of 7.40 mM MgCl2
Answer #1

Sol.

As The Conc. of MgCl2 = C = 7.40 mM = 7.40 × 10-3 M

Dissociation reaction of MgCl2 :

MgCl2 ----> Mg2+ + 2Cl-

So , Van't Hoff factor = i = Total Number of ions produced

= 1 + 2 = 3  

Gas constant = R = 0.0821 L atm / K mol  

Temeprature = T  

So , Osmotic Pressure  = P = iCRT

= 3 × 7.40 × 10-3 × 0.0821 × T

= 0.00182262 × T

As the temperature , T( in K )  is not given in the question , so , i am writing it as only T